OpenStudy (anonymous):

Fool's problem of the day, Find the sum of \( n \) terms of the series: \[ \frac 13+ \frac{2}{21}+ \frac{3}{91} + \frac{4}{273} \cdots \text{ upto n terms } \]

OpenStudy (lgbasallote):

is this arithmetic progression?

OpenStudy (anonymous):

Noooway, this isn't AP.

OpenStudy (lgbasallote):

fool outdid himself today :P i cant find a pattern or im just that weak hahaha

OpenStudy (anonymous):

This took me about 10 mints to solve (without electronic aid), lets see how much you guys take :)

OpenStudy (anonymous):

And please no cheating :P

OpenStudy (lgbasallote):

it has been 8 minutes so far.....

OpenStudy (lgbasallote):

haha finally found the pattern took me long enough :))

OpenStudy (anonymous):

Yes sure you used google ? :P

OpenStudy (lgbasallote):

no...i looked at the last term :P but i still dunno the sum

OpenStudy (anonymous):

It's the pattern which is a bit elusive, the rest is kids stuff! ;)

OpenStudy (callisto):

the denominator is like, 1x3, 3x7, 7x13, 13x21 and the numerator is like the difference of the 2 numbers divided by 2.. Nah.. but don't even know what to do..

OpenStudy (lgbasallote):

there's no pattern in the denominator..it is inconsistent in callisto's interpretation i think...the pairs are 1-3, 3-7, 7-13, 13-21...there is no pattern on the second digit of the pair

OpenStudy (lgbasallote):

oh wait there was...i stand corrected

OpenStudy (callisto):

anyway, i was incorrect then :D

OpenStudy (lgbasallote):

so next one is gonna be 5/ 756?

OpenStudy (anonymous):

No, the next one is not 756.

OpenStudy (lgbasallote):

651?

OpenStudy (anonymous):

HINT: The denominator is \( 1+n^2+n^4 \)

OpenStudy (lgbasallote):

where n = numerator lol

OpenStudy (lgbasallote):

oh wait yeah..

OpenStudy (anonymous):

I would leave the rest to your envisage.

OpenStudy (anonymous):

How did you figure out that FoolForMath? "\(1 + n^2 + n^4\)"

OpenStudy (lgbasallote):

so the form is n/(n^4 + n^2 + 1)

OpenStudy (anonymous):

Hmm, I did something like this before Ishaan ;)

OpenStudy (mani_jha):

\\[n/(n ^{4}+2n ^{2}+1)-n ^{2}=n/(n ^{2}+1)^{2}-n ^{2}\]

OpenStudy (anonymous):

Nice problem though, most of the users would fail to see such pattern.

OpenStudy (anonymous):

The problem is not finished yet. The rest is somewhat interesting :)

OpenStudy (mani_jha):

Do we need to use the telescoping method?

OpenStudy (anonymous):

Yes, that's a way to do the rest, but we need to make it ready first :)

OpenStudy (mani_jha):

Oh well let me think..

OpenStudy (lgbasallote):

you have to brew the ingredients first

OpenStudy (mani_jha):

We've to express the general term(nth term) as a difference of the nth and (n+1)th terms, or nth and (n-1)th terms. Come on guys and gals, think!

OpenStudy (anonymous):

\[\frac{n}{(n^2 + 1 - n)(n^2+1+n)} = \frac 1 2\left(\frac{1}{n^2+1 - n} - \frac{1}{n^2 +1+n}\right)\]

OpenStudy (mani_jha):

\[\sum_{1}^{n}n/(n^4+n^2+1)=\sum_{1}^{n}(n^2+n+1)-(n^2-n+1)/2(n^2+n+1)(n^2-n+1)\] =\[\sum_{1}^{n}((1/n^2-n+1)-(1/n^2+n+1))/2\]

OpenStudy (mani_jha):

Oh, it's already done. Now just substituting n=1, n=2 etc would do

OpenStudy (anonymous):

Is it 1?

OpenStudy (mani_jha):

Yes, it seems so. If we put n=1 we get 1 for the first term and the other terms keep getting cancelled on summation.

OpenStudy (anonymous):

A strange answer for such a complex series :/

OpenStudy (callisto):

\[\frac{n}{(n^2 + 1 - n)(n^2+1+n)} = \frac 1 2\left(\frac{1}{n^2+1 - n} - \frac{1}{n^2 +1+n}\right)\] Is it something related to partial fraction?

OpenStudy (anonymous):

Yeah, kind of

OpenStudy (lgbasallote):

looooong equaion @_@ too many numbers...fainting

OpenStudy (callisto):

no wonder why i don't understand.. thanks ..

OpenStudy (mani_jha):

That's not so difficult, Callisto. It is similar to the following operation: \[1/n(n+1)=(n+1)-n/n(n+1)\] =\[(n+1)/n(n+1)-n/n(n+1)=1/n-1/(n+1)\] If you can understand the above, you would've no problem understanding this: He just wrote \[n=2n/2={(n^2+n+1)-(n^2-n+1)}/2\] Then each term above was divided by the denominator. Ok?

OpenStudy (lgbasallote):

haha 1 hour...it seems Fool is smarter than any of us or all of us combined :P

OpenStudy (anonymous):

yeahh :/

OpenStudy (callisto):

but sometimes it's difficult to think how to break down the fraction, my teacher hasn't taught us anything about partial fraction except 1/[n(n+1)] = 1/n - 1/(n+1).. (just remembering) So, whenever i see something related to partial fraction, probably i'll give up at this stage :S

OpenStudy (lgbasallote):

you'll learn it soon :P

OpenStudy (callisto):

After seeing similar questions for thousand times, I will :D

OpenStudy (lgbasallote):

there are actually cases of partial fractions

OpenStudy (callisto):

So the number of ways to do it is limited?